# Learning with Local Search MCMC Layers

### v1

@@ C.4 Proof of strict convexity
$\leq\frac{2R_{\mathcal{C}}}{t}||\bm{\theta}||+\frac{2R_{\varphi}}{t},$
where $R_{\mathcal{C}}=\max_{\bm{y}\in\mathcal{Y}}||\bm{y}||$ and $R_{\varphi}=\max_{\bm{y}\in\mathcal{Y}}|\varphi(\bm{y})|$. Thus, we have:
$\lambda^{\prime}_{2}\leq 1-G|\mathcal{Y}|\exp\left(-\frac{4R_{\varphi}}{t}\right)\exp\left(-\frac{4R_{\mathcal{C}}}{t}||\bm{\theta}||\right),$
@@
$\lambda_{\bm{\theta}}\leq 1-\frac{G|\mathcal{Y}|\exp\left(-\frac{4R_{\varphi}}{t}\right)}{2}\exp\left(-\frac{4R_{\mathcal{C}}}{t}||\bm{\theta}||\right),$
@@
C.7 Proofs of Section 4.3 and Section B.1
Section 4.3.
The distribution of the first iterate of the Markov chain with transition kernel defined in Eq. 3 and initialized at the ground-truth structure ${\bm{y}}$ is given by:
@@
=$=\begin{cases}q({\bm{y}},\bm{y}^{\prime})\min\left[1,\frac{q(\bm{y}^{\prime},{\bm{y}})}{q({\bm{y}},\bm{y}^{\prime})}\exp\left(\left[\langle\bm{\theta},\bm{y}^{\prime}-{\bm{y}}\rangle+\varphi(\bm{y}^{\prime})-\varphi({\bm{y}})\right]/t\right)\right]&\text{if $\bm{y}^{\prime}\in\mathcal{N}({\bm{y}})$,}\\ 1-\sum_{\bm{y}^{\prime\prime}\in\mathcal{N}({\bm{y}})}(\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}})(\bm{y}^{\prime\prime})&\text{if $\bm{y}^{\prime}={\bm{y}}$,}\\ 0&\text{else.}\end{cases}$
@@
Moreover, we have that $f_{{\bm{y}}}(\,\cdot\,;\bm{y}^{\prime})$ is actually differentiable everywhere as its gradient can be continuously extended to the junction affine hyperplane with constant value equal to $q({\bm{y}},\bm{y}^{\prime})(\bm{y}^{\prime}-{\bm{y}})$. We now show that $f_{{\bm{y}}}(\,\cdot\,;\bm{y}^{\prime})$ is $\frac{1}{t}q({\bm{y}},\bm{y}^{\prime})\cdot||\bm{y}^{\prime}-{\bm{y}}||^{2}$-smooth. Indeed, it is defined as the composition of the linear form $\bm{\theta}\mapsto\langle\bm{\theta},\bm{y}^{\prime}-{\bm{y}}\rangle$ and the function $g:\mathbb{R}\rightarrow\mathbb{R}$ given by:
@@
$||\nabla_{\bm{\theta}}f_{{\bm{y}}}(\bm{\theta_{1}},\bm{y}^{\prime})-\nabla_{\bm{\theta}}f_{{\bm{y}}}(\bm{\theta_{2}},\bm{y}^{\prime})||$ $=||g^{\prime}(\langle\bm{\theta}_{1},\bm{y}^{\prime}-{\bm{y}}\rangle)(\bm{y}^{\prime}-{\bm{y}})-g^{\prime}(\langle\bm{\theta}_{2},\bm{y}^{\prime}-{\bm{y}}\rangle)(\bm{y}^{\prime}-{\bm{y}})||$
$=|g^{\prime}(\langle\bm{\theta}_{1},\bm{y}^{\prime}-{\bm{y}}\rangle)-g^{\prime}(\langle\bm{\theta}_{2},\bm{y}^{\prime}-{\bm{y}}\rangle)|\cdot||\bm{y}^{\prime}-{\bm{y}}||$
$\leq\frac{1}{t}q({\bm{y}},\bm{y}^{\prime})\cdot|\langle\bm{\theta}_{1},\bm{y}^{\prime}-{\bm{y}}\rangle-\langle\bm{\theta}_{2},\bm{y}^{\prime}-{\bm{y}}\rangle|\cdot||\bm{y}^{\prime}-{\bm{y}}||$
$\leq\frac{1}{t}q({\bm{y}},\bm{y}^{\prime})\cdot||\bm{y}^{\prime}-{\bm{y}}||^{2}\cdot||\bm{\theta}_{1}-\bm{\theta}_{2}||,$
and $f_{{\bm{y}}}(\,\cdot\,,\bm{y}^{\prime})$ is $\frac{1}{t}q({\bm{y}},\bm{y}^{\prime})\cdot||\bm{y}^{\prime}-{\bm{y}}||^{2}$-smooth. Thus, recalling that $F_{{\bm{y}}}$ is defined as
@@
we have that $F_{{\bm{y}}}$ is $\sum_{\bm{y}^{\prime}\in\mathcal{N}({\bm{y}})}\frac{1}{t}q({\bm{y}},\bm{y}^{\prime})\cdot||\bm{y}^{\prime}-{\bm{y}}||^{2}=\mathbb{E}_{q({\bm{y}},\,\cdot\,)}||Y-{\bm{y}}||_{2}^{2}/t$-smooth. Finally, as $\Omega_{{\bm{y}}}\coloneqq(F_{{\bm{y}}})^{*}$, Fenchel duality theory gives that $\Omega_{{\bm{y}}}$ is $t/\mathbb{E}_{q({\bm{y}},\,\cdot\,)}||Y-{\bm{y}}||_{2}^{2}$-strongly convex.
@@
$\forall\bm{\theta}\in\mathbb{R}^{d},\,||\nabla F_{{\bm{y}}}(\bm{\theta})||\leq R_{\mathcal{N}({\bm{y}})}\coloneqq\max_{\bm{\mu}\in\operatorname{\mathsf{conv}}\left(\mathcal{N}({\bm{y}})\cup\{{\bm{y}}\}\right)}||\bm{\mu}||,$
@@
Thus, the $\mathbb{E}_{q({\bm{y}},\,\cdot\,)}||Y-{\bm{y}}||_{2}^{2}/t$-smoothness of $L_{\Omega_{{\bm{y}}}}(\,\cdot\,;{\bm{y}})$ follows directly from the previously established $\mathbb{E}_{q({\bm{y}},\,\cdot\,)}||Y-{\bm{y}}||_{2}^{2}/t$-smoothness of $F_{{\bm{y}}}$. Similarly, the expression of $\nabla_{\bm{\theta}}L_{\Omega_{{\bm{y}}}}(\bm{\theta}\,;{\bm{y}})$ follows from the previously established expression of $\nabla_{\bm{\theta}}F_{{\bm{y}}}(\bm{\theta})$, and we have:
@@
Section B.1.
In the unsupervised setting, given a dataset $({\bm{y}}_{i})_{i=1}^{N}$, the distribution of the first iterate of the Markov chain with transition kernel defined in Eq. 3 and initialized by ${\bm{y}}^{(0)}={\bm{y}}_{i}$, with $i\sim\mathcal{U}(\llbracket 1,N\rrbracket)$, is given by:
@@
Let $F_{\bar{Y}_{N}}\coloneqq\frac{1}{N}\sum_{i=1}^{N}F_{{\bm{y}}_{i}}$ Then, the exact same arguments as in the supervised case hold, and the results of Section B.1 are obtained by replacing $F_{{\bm{y}}}$ by $F_{\bar{Y}_{N}}$ in the proof of Section 4.3, and noticing that the previously shown ${\mathbb{E}}_{q({\bm{y}}_{i},\,\cdot\,)}||Y-{\bm{y}}_{i}||^{2}_{2}/t$-smoothness of $F_{{\bm{y}}_{i}}$ gives that $F_{\bar{Y}_{N}}$ is $\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}_{q({\bm{y}}_{i},\,\cdot\,)}||Y-{\bm{y}}_{i}||^{2}_{2}/t$-smooth. Similar arguments also hold for the regularized optimization formulation, by noting that this time we have $\nabla F_{\bar{Y}_{N}}({\mathbb{R}}^{d})\subset\operatorname{\mathsf{conv}}\left(\bigcup_{i=1}^{N}\left\{\mathcal{N}({\bm{y}}_{i})\cup\{{\bm{y}}_{i}\}\right\}\right)$. ∎

### v2

@@ D.4 Proof of strict convexity
[…]
@@
[…]
@@
D.7 Proofs of Section˜4.3 and Section˜B.1
Section˜4.3.
The distribution of the first iterate of the Markov chain with transition kernel defined in Eq.˜3 and initialized at the ground-truth structure $\bm{y}$} is given by:
@@
$=\begin{cases}q({\bm{y}},\bm{y}^{\prime})\min\left[1,\frac{q(\bm{y}^{\prime},{\bm{y}})}{q({\bm{y}},\bm{y}^{\prime})}\exp\left(\left[\langle\bm{\theta},\bm{y}^{\prime}-{\bm{y}}\rangle+\varphi(\bm{y}^{\prime})-\varphi({\bm{y}})\right]/t\right)\right]&\text{if $\bm{y}^{\prime}\in\mathcal{N}({\bm{y}})$,}\\ 1-\sum_{\bm{y}^{\prime\prime}\in\mathcal{N}({\bm{y}})}(\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}})(\bm{y}^{\prime\prime})&\text{if $\bm{y}^{\prime}={\bm{y}}$,}\\ 0&\text{else.}\end{cases}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Section˜B.1.
In the unconditional setting, given a dataset $({\bm{y}}_{i})_{i=1}^{N}$, the distribution of the first iterate of the Markov chain with transition kernel defined in Eq.˜3 and initialized by ${\bm{y}}^{(0)}={\bm{y}}_{i}$, with $i\sim\mathcal{U}(\llbracket 1,N\rrbracket)$, is given by:
@@
[…]
D.8 Proof of Section˜B.2
Proof.
The first point is directly given by the fact that ${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]$ is the expectation of a distribution over ${\mathcal{N}}({\bm{y}})\cup\{{\bm{y}}\}$. For the second and third points, as derived in Section˜D.7, we have:
${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]={\bm{y}}+\!\!\!\!\!\!\sum_{\bm{y}^{\prime}\in\mathcal{N}_{{\bm{y}}}^{-}(\bm{\theta})}\!\!q(\bm{y}^{\prime},{\bm{y}})\exp\left(\left[\langle\bm{\theta},\bm{y}^{\prime}-{\bm{y}}\rangle+\varphi(\bm{y}^{\prime})-\varphi({\bm{y}})\right]/t\right)\cdot\left(\bm{y}^{\prime}-{\bm{y}}\right)\;+\!\!\!\!\!\sum_{\bm{y}^{\prime}\in\mathcal{N}_{{\bm{y}}}^{+}(\bm{\theta})}\!\!q({\bm{y}},\bm{y}^{\prime})\cdot\left(\bm{y}^{\prime}-{\bm{y}}\right).$
Define then:
${\mathcal{N}}_{\text{better}}({\bm{y}})$ $\coloneqq\left\{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})\mid\langle{\bm{\theta}},{\bm{y}}^{\prime}\rangle+\varphi({\bm{y}}^{\prime})>\langle{\bm{\theta}},{\bm{y}}\rangle+\varphi({\bm{y}})\right\},$
${\mathcal{N}}_{\text{worse}}({\bm{y}})$ $\coloneqq\left\{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})\mid\langle{\bm{\theta}},{\bm{y}}^{\prime}\rangle+\varphi({\bm{y}}^{\prime})<\langle{\bm{\theta}},{\bm{y}}\rangle+\varphi({\bm{y}})\right\}$
as the sets of improving and worsening neighbors of $\bm{y}$} respectively (assuming no neighbor of $\bm{y}$} has exactly equal objective value for simplicity, which is true almost everywhere w.r.t. ${\bm{\theta}}\in{\mathbb{R}}^{d}$).
Low temperature limit.
We have:
${\mathcal{N}}_{{\bm{y}}}^{+}({\bm{\theta}})\xrightarrow[t\rightarrow 0^{+}]{}{\mathcal{N}}_{\text{better}}({\bm{y}}),\quad\text{and}\quad{\mathcal{N}}_{{\bm{y}}}^{-}({\bm{\theta}})\xrightarrow[t\rightarrow 0^{+}]{}{\mathcal{N}}_{\text{worse}}({\bm{y}}).$
Then, as $x<0\implies\exp(x/t)\xrightarrow[t\rightarrow 0^{+}]{}0$, we have effectively
${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]\xrightarrow[t\rightarrow 0^{+}]{}{\bm{y}}+\sum_{{\bm{y}}^{\prime}\in{\mathcal{N}}_{\text{better}}({\bm{y}})}q({\bm{y}},{\bm{y}}^{\prime})\cdot({\bm{y}}^{\prime}-{\bm{y}}).$
High temperature limit.
As $\forall x\in{\mathbb{R}},\;\exp(x/t)\xrightarrow[t\rightarrow\infty]{}1$, we have:
${\mathcal{N}}_{{\bm{y}}}^{+}({\bm{\theta}})\xrightarrow[t\rightarrow\infty]{}\left\{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})\mid q({\bm{y}}^{\prime},{\bm{y}})>q({\bm{y}},{\bm{y}}^{\prime})\right\},\quad\text{and}\quad{\mathcal{N}}_{{\bm{y}}}^{-}({\bm{\theta}})\xrightarrow[t\rightarrow\infty]{}\left\{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})\mid q({\bm{y}}^{\prime},{\bm{y}})\leq({\bm{y}},{\bm{y}}^{\prime})\right\}.$
Thus, we have:
${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]\xrightarrow[t\rightarrow\infty]{}{\bm{y}}+\sum_{{\bm{y}}^{\prime}\mid q({\bm{y}}^{\prime},{\bm{y}})\leq({\bm{y}},{\bm{y}}^{\prime})}q({\bm{y}}^{\prime},{\bm{y}})\cdot({\bm{y}}^{\prime}-{\bm{y}})+\sum_{{\bm{y}}^{\prime}\mid q({\bm{y}}^{\prime},{\bm{y}})>({\bm{y}},{\bm{y}}^{\prime})}q({\bm{y}},{\bm{y}}^{\prime})\cdot({\bm{y}}^{\prime}-{\bm{y}}),$
which gives effectively:
${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]\xrightarrow[t\rightarrow\infty]{}{\bm{y}}+\sum_{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})}\min\left[q({\bm{y}},{\bm{y}}^{\prime}),q({\bm{y}}^{\prime},{\bm{y}})\right]\cdot({\bm{y}}^{\prime}-{\bm{y}}).$